OpenStudy (anonymous):

james weight decreased by 5% after dieting. he now weighs 133 pounds. what was his original weight?

OpenStudy (anonymous):

see below,, 1st find 1% of 133 as 133/100=1.33 now for 5 % multiply it by 5i.e. 5*1.3=6.5 now add this up to 133 i.e. 133+6.5=139.5

OpenStudy (dls):

\[\Large x-\frac{5}{100}x =133\] solve for x

OpenStudy (driftracer305):

@zaibali.qasmi ......... it is 139.65....to be exact

OpenStudy (kropot72):

133 pounds is 0.95 of his original weight. Original weight is 133/0.95 = 140 pounds

OpenStudy (dls):

140 kgs nearly

OpenStudy (kropot72):

140 pounds exactly.

OpenStudy (anonymous):

1st find 1% of 133 as 133/100=1.33 now for 5 % multiply it by 5i.e. 5*1.3=6.65 now add this up to 133 i.e. 133+6.5=139.65

OpenStudy (driftracer305):

yup.....@zaibali.qasmi ........u got it dude

OpenStudy (anonymous):

Thanks everyone I get it :)

OpenStudy (anonymous):

thanxx for correction.. @driftracer305

OpenStudy (anonymous):

u are warmly welcome @chris117

OpenStudy (driftracer305):

hey..........lol @zaibali.qasmi ............sorry dude.....didnt mean to be rude thou......m sure ur far way better at it than me.....

OpenStudy (anonymous):

Youre welcome chris117